OpenStudy (anonymous):

solve for all values of x: 2x^3-3x^2-8x+12=0 are

OpenStudy (anonymous):

I'll give you the first step and you will do it, OK?

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

2x^3-3x^2-8x+12=0 write it in a different order 2x^3-8x-3x^2+12=0 factor 2x(x^2-4)-3(x^2-4)=0 (2x-3)(x^2-4)=0 Go

OpenStudy (anonymous):

(HINT: one of the parenthesis has to be equal zero)

OpenStudy (anonymous):

So (2x-3)=0 and (x^2-4)=0 2x-3=0 x^2-4=0 x=? x=? Solve for x in each one....

OpenStudy (anonymous):

thank you! appreciate the assistance.

OpenStudy (anonymous):

Anytime!

OpenStudy (anonymous):

I don't know what x equals on the 2x-3=0 side.. confused

OpenStudy (anonymous):

2x-3=0 x^2-4=0 2x-3 + 3 = 0 + 3 x^2-4 + 4 =0 + 4 2x = 3 x^2 = 4 x= 3/2 x= 2, -2

OpenStudy (anonymous):

tnx for the medal brah!

OpenStudy (anonymous):

thanks 4 your help. I understand now. and your welcome!!

OpenStudy (anonymous):

Anytime!

OpenStudy (anonymous):

I know what it feels like when you don't get your questions answered and explained. I had this in Literature, and that's why i helped you.